OpenStudy (anonymous):

Write the equation of direct variation for x = 4, y = -2.1276.

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

help?

OpenStudy (skullpatrol):

Do you know the form of a "direct variation"?

OpenStudy (anonymous):

y=kx?

OpenStudy (skullpatrol):

Correct.

OpenStudy (anonymous):

what now?

OpenStudy (skullpatrol):

The question gives you x = 4, y = -2.1276 what do you think?

OpenStudy (anonymous):

-2.12176=k4?

OpenStudy (skullpatrol):

Perfect, now what is the only "unknown" in that equation?

OpenStudy (anonymous):

K?

OpenStudy (skullpatrol):

Yes....

OpenStudy (anonymous):

what do i do?

OpenStudy (anonymous):

idk

OpenStudy (skullpatrol):

$$\Huge-2.12176=k*4$$

OpenStudy (skullpatrol):

Once you know what "k" is you can put it into the direct variation form y = kx, right?

OpenStudy (anonymous):

k=-.53044?

OpenStudy (anonymous):

i think?

OpenStudy (skullpatrol):

Yes, that is what $$\frac{-2.12176}{4}$$equals :)

OpenStudy (anonymous):

ok thankss

OpenStudy (skullpatrol):

Now answer the question and put into the y=kx form please.

OpenStudy (skullpatrol):

Write the equation of direct variation

OpenStudy (anonymous):

-2.12176=-.53044(4)

OpenStudy (anonymous):

??

OpenStudy (skullpatrol):

That equation says: -2.12176 = -.53044(4) -2.12176 = -2.12176 which is a true statement. The question asks for "Write the equation of direct variation"

OpenStudy (anonymous):

so what do i do?

OpenStudy (skullpatrol):

Substitute the value of "k" you calculated into the general form y = kx.

OpenStudy (anonymous):

so my answer is -.53044?

OpenStudy (skullpatrol):

Your answer is the asked for equation.

OpenStudy (skullpatrol):

k = -0.53044

OpenStudy (anonymous):

thank youuuu so much

OpenStudy (skullpatrol):

$$\Huge y= -0.53044x$$

OpenStudy (skullpatrol):

Since y = kx

OpenStudy (anonymous):

thanks brah

OpenStudy (skullpatrol):

Thanks for trying to learn.